### v1

@@ 3. Robustly Periodic chain-recurrence classes
Before proceeding to step 2 of our argument, we present the next result about the non-trivial dynamical behavior of the chain-recurrence classes $C(\sigma_{Y})$ sufficiently $C^{1}$-close to $X\in\mathcal{R}$.
Lemma 3.8.
Let $X\in\mathcal{R}$ and let $C(\sigma_{X})\subset M$ be a non-trivial sectional-hyperbolic chain-recurrent class for $X$. There is a neighborhood $\mathcal{U}_{2}\subset\mathcal{U}_{1}$ of $X$ such that for every $Y\in\mathcal{U}_{2}$ one has $C(\sigma_{Y})\not\subset W^{s}(Sing(Y)$.
Proof.
By lemma 3.5 we have that $C(\sigma_{X})$ is a homoclinic class. Take two different periodic orbits $\gamma_{0},\gamma_{1}$ in $C(\sigma_{X})$, and let $d_{\gamma}=d_{H}(\gamma_{0},\gamma_{1})>0$. By continuity of $Y\in\mathcal{U}_{X}\to C(\gamma_{Y})$ there is a neighborhood $\mathcal{U}_{2}\subset\mathcal{U}_{1}$ of $X$ such that
@@
Let $X\in\mathcal{R}$ and let $C(\sigma_{X})\subset M$ be a non-trivial sectional-hyperbolic chain-recurrent class for $X$. There is a neighborhood $\mathcal{U}\subset\mathcal{U}_{2}$ of $X$ such that, for every $Y\in\mathcal{U}$, the periodic orbit $\gamma_{Y}$ given by Lemma 3.6 is chain-attainable from $\sigma_{Y}$.
@@
To begin with, we recall that by Lemma 2.2 and Lemma 3.5, $C(\sigma_{X})$ is a homoclinic class whose singularities are Lorenz-like. Let us denote
@@
Notice that for any vector field $Y$ sufficiently $C^{1}$-close to $X$, the continuations $Sing(X|_{C(\sigma_{Y})})$ consist entirely of Lorenz-like singularities.
@@
Also, from the proof of Lemma 3.5, there is a sequence of periodic orbits $\gamma_{n}\subset C(\sigma_{X})$ such that $d_{H}(\gamma_{n},C(\sigma_{X}))\to 0$. In this case, let us assume
@@
for every $n\geq 0$. In particular, for each i=x,1,2,3,,ki=x,1,2,3,...,k, we can obtain sequences of points $(p_{n,i}^{s})_{n\geq 1}$ and $(p_{n,i}^{u})_{n\geq 1}$ such that for every i=0,1,2,,ki=0,1,2,...,k:
@@
$\mathcal{N}_{i}^{in}=\mathcal{N}^{in}_{p_{s,i}}\subset B(\sigma_{i},\delta)\textrm{ and }\mathcal{N}_{i}^{out}=\mathcal{N}^{out}_{p_{u,i}}\subset B(\sigma_{i},\delta)$
@@
By the continuity of $Y\in\mathcal{U}_{X}\to C(\gamma_{Y})$, there is a $C^{1}$-neighborhood $\mathcal{U}\subset\mathcal{U}_{2}$ of $X$ such that
@@
Now, we are ready to proceed with the proof. Let us split it into two cases.
Case 1: There is $x\in C(\sigma_{Y})$ such that $\omega(x)\cap Sing(Y)=\emptyset$.
Since $\omega(x)$ has no singularities, it follows by Lemma 2.1 that it is hyperbolic of saddle type. In particular, has also local stable and unstable manifolds of size $\varepsilon>0$. So, by shrinking $\mathcal{U}$ if it is necessary, we obtain that
$W^{u}(z)\pitchfork W^{s}(\gamma_{Y})\neq\emptyset$
for some $z\in\omega(x)\subset C(\sigma_{Y})$. Thus, since $z$ is chain attainable from $\sigma_{Y}$, one has that $\gamma_{Y}$ is chain attainable from $\sigma_{Y}$.
Case 2: $\omega(x)\cap Sing(Y)\neq\emptyset$ for any $x\in C(\sigma_{Y})$.
In this case, we denote
The compactness of $C(\sigma_{Y})\setminus B(Sing(Y|_{C(\sigma_{Y})},\delta)$ implies the existence of a positive number $T_{\delta}$ such that if
$z\in C(\sigma_{Y})\setminus B(Sing(Y|_{C(\sigma_{Y})},\delta),$
there is $0<t_{z}\leq T_{\delta}$ such that $Y_{t_{z}}(z)\in\mathcal{N}^{in}$ and $Y_{(0,t_{z})}(z)\cap\mathcal{N}^{in}=\emptyset$.
By Lemma 3.8, there is a point $x\in C(\sigma_{Y})\setminus W^{s}(Sing(Y))$. Moreover, since the positive orbit of $x$ accumulates $Sing(Y|_{C(\sigma_{Y})})$, there is a sequence of times $t_{n}\to\infty$ and a singularity $\sigma^{i}\in Sing(Y|_{C(Y)})$ such that $X_{t_{n}}(x)\to\sigma^{i}$. In particular, the positive orbit of $x$ crosses $\mathcal{N}^{in}$ infinitely many times. We may assume $x\in\mathcal{N}^{in}_{i}$ for some $i\in\{0,1,\ldots,k\}$. Let us consider a box $B\subset\mathcal{N}^{in}_{i}$ containing $x$ with central volume $0<b<\frac{d^{s}_{Y}}{2}$.
Claim: For any $T>0$, there is $T_{b}>T$, depending on $b$, such that $Y_{[0,T_{b}]}(B)\pitchfork W^{s}(\gamma_{Y})\neq\emptyset$.
$K=\inf\left\{\frac{\|Y(z)\|}{\|Y(w)\|}:z,w\notin B(\sigma_{Y},\delta)\right\},$
the sectional-hyperbolicity implies
(5) $)\displaystyle\|\pi_{N_{Y_{t}(x)}}(\Phi_{t}(\hat{v}$$)\|$ $=$ $\frac{\|Y(x)\|}{\|Y(Y_{t}(x))\|}\cdot\det(\Phi^{Y}_{t}(x)|_{L_{x}})$
$\geq$ $\frac{\|Y(x)\|}{\|Y(Y_{t}(x))\|}\cdot e^{\lambda t}$
$\geq$ $Ke^{\lambda t},$
whenever $t>0$, $Y_{[0,t)}(x)\in U$ and $Y_{t}(x)\notin B(\sigma,\delta)$.
Recall by the construction of the cross sections $\mathcal{N}^{in}$ and $\mathcal{N}^{out}$ that if $z\in\mathcal{N}^{in}_{i}$, then $Y_{[0,\tau(z))}(z)\in B(\sigma_{Y},\delta)\subset U$ and $H_{i}^{Y}(z)=Y_{\tau(z)}(z)\in\mathcal{N}^{out}_{i}$. In this case, define
$\tau_{i}=\inf\{\tau(z):z\in\mathcal{N}^{in}_{i}\}.$
Notice that we can choose $\delta>0$ in a such way that
where $0<\lambda_{u}<\lambda$. Let $\mathcal{N}_{X_{t}(x)}$, $t>\tau(x)$, be a cross-section through $X_{t}(x)$ of radius at least $\eta_{\delta}$. Since $C(\sigma_{Y})$ is not necessarily Lyapunov stable, the set $H^{t}_{x}(B)$ may not be entirely contained in $B(C(\sigma_{Y}),2\delta)$ for some $t>\tau_{i}$. Nevertheless, there is a connected component $B_{1}$ of $H_{x}^{t}(B)$ whose central volume is bigger than $b$ and $\overline{B}\subset B(C(\sigma_{Y}),2\delta)$. In this case, we set
$r_{1}=\min\{r>0:Y_{r}(z)\in\partial B(C(\sigma_{Y}),2\delta)\text{ for some }z\in(H^{r}_{x})^{-1}(B_{1})\}.$
Hence, the map
satisfies
$H_{i}^{Y}(B)\subset Dom(H^{r1}_{x})\textrm{ and }|H^{t}_{x}(B)|_{c}=|H_{x}^{r_{1}}\circ H_{i}(B)|_{c}\geq e^{\lambda^{\prime}(\tau_{i}+r)}b.$
In particular, one has
$|H_{x}^{t}(B)|_{c}\geq e^{\tau_{i}+r_{1}}b>2b.$
By the choice of $\ell$, there is $s\geq 0$ such that $\tau+r_{1}+\cdots+r_{\ell}+s=t^{1}_{x}$. Besides, if $)\hat{w}=\pi_{N_{Y_{t_{x}^{1}-s}(x)}}(\Phi_{t^{1}_{x}-s}(\hat{v}$)), we have by (5) and (1) that
$\|\pi_{N_{Y_{t_{x}^{1}}(x)}}(\Phi_{s}(\hat{w}))\|\geq\frac{\|Y(x)\|}{\|Y(Y_{t}(x))\|}\cdot e^{\lambda s}\geq e^{\lambda s}>1.$
By repeating the previous process in an inductively way, there is $N\in\mathbb{N}$, and a connected component $C_{N}\subset\mathcal{N}^{in}$ of $H_{x}^{t_{x}^{N}}(B)$ whose central volume is bigger than $Nb>\frac{3}{2}d_{Y}^{s}$. Furthermore, we can choose $N$ and $C_{N}$ in a such way that $t_{x}^{N}>T$ and $C_{N}\subset\mathcal{N}_{i}^{in}$. So, we have that $Y_{[0,s]}(C_{N})$, $s>0$, is a submanifold of dimension $\dim F$ which intersects $W^{s}(\sigma_{j})$ transversely for some $j\in\{0,1,\ldots,k\}$. Moreover, since the central volume of $C_{N}$ is bigger than $d^{s}_{Y}$, one has by (b) that $W^{s}(\gamma_{j})\pitchfork Y_{[0,s_{0}]}(C_{N})\neq\emptyset$ for some $s_{0}>0$. So, the claim follows by taking $T_{b}=t_{x}^{N}+s_{0}$.
To finish the proof, recall that the positive orbit of $x$ crosses $\mathcal{N}^{in}$ infinitely many times and, therefore, there is a sequence of times $t_{n}\to\infty$ and a singularity $\sigma^{i}\in Sing(Y|_{C(Y)})$ such that $X_{t_{n}}(x)\to x_{0}$ and $x_{0}\in W^{s}_{loc}(\sigma^{i})$. Without loss of generality, we can assume $x_{0},X_{t_{n}}(x)\in\mathcal{N}^{out}_{i}$, for every $n>0$. Fix $\varepsilon$, $T>0$ and fix $N>0$ such that $d(X_{T_{n}}(x),x_{0})\leq\varepsilon$, for every $n\geq N$. Fix $s_{\varepsilon}>T$ satisfying $d(X_{-s_{\varepsilon}}(x_{0}),\sigma^{i})\leq\varepsilon$. Let $x_{n}\in\mathcal{N}^{i}n$ be such that $x_{n}=X_{t_{N}+s}(x)$ and $s>T$. We we consider a box $B$ around $x_{n}$ with diameter $\varepsilon$ and volume $b$, by the claim and the invariance of stable sets, there is $z\in B\cap W^{s}(\gamma_{Y})$. Therefore, there exist $s_{z}>T$ and $x_{z}\in\gamma_{Y}$ such that $d(X_{s_{z}}(z),x_{z})\leq\varepsilon$. Finally, the set
$\{\sigma^{i},X_{-s_{\varepsilon}}(x_{0}),X_{t_{n}}(x),z,x_{z}\}$
is an $\varepsilon$-$T$-chain from $\sigma^{i}$ to $\gamma_{Y}$. Since $\sigma_{Y}$ and $\sigma^{i}$ are chain related, the proof is complete. ∎
Next, we are ready to prove Theorem 3.1.
Proof of Theorem 3.1.
Let $\mathcal{R}$ be given by Lemma 3.2. First, by Lemma 3.5 we have that $C(\sigma_{X})$ is a homoclinic class, which implies by Lemma 3.8 that the continuation $C(\sigma_{Y})$ does not consist of a homoclinic loop around $\sigma_{Y}$ for every $Y\in\mathcal{U}$. Second, by Lemma 3.6 and Lemma 3.9, there is a periodic orbit $\gamma_{Y}$ such that $\sigma_{Y}\sim\gamma_{Y}$ and $\gamma_{Y}\sim\sigma_{Y}$, showing that $\gamma_{Y}\in C(\sigma_{Y})$ by definition of chain class. ∎
4. Proof of Theorems A and B

### v2

@@ 3. Robustly Periodic chain-recurrence classes
Let $X\in\mathcal{R}$ and let $C(\sigma_{X})\subset M$ be a non-trivial sectional-hyperbolic chain-recurrent class for $X$, which is not reduced to either a homoclinic loop or a saddle connection between singularities. There is a neighborhood $\mathcal{U}\subset\mathcal{U}_{2}$ of $X$ such that, for every $Y\in\mathcal{U}$, the periodic orbit $\gamma_{Y}$ given by Lemma 3.12 is chain-attainable from $\sigma_{Y}$.
@@
First, we recall that by Lemma 2.2 the singularities of $C(\sigma_{X})$ are Lorenz-like. Let us denote
@@
Also, from item (2) of Lemma 3.2 there is a sequence of periodic orbits $\{\gamma_{n}\}_{n\geq 1}$ of $X$ such that $d_{H}(\gamma_{n},C(\sigma_{X}))\to 0$. In this case, let us assume
@@
for every $n\geq 0$. In particular, for each i=0,,,ki=0,,...,k, we can obtain sequences of points $(p_{n,i}^{s})_{n\geq 1}$ and $(p_{n,i}^{u})_{n\geq 1}$ such that for every i=0,,ki=0,...,k:
@@
$\mathcal{N}_{i}^{in}=\mathcal{N}^{in}_{p_{s,i}}\subset\partial B(\sigma_{i},\delta)\textrm{ and }\mathcal{N}_{i}^{out}=\mathcal{N}^{out}_{p_{u,i}}\subset\partial B(\sigma_{i},\delta)$
@@
Now, let consider the following cases:
Case 1: There is $x\in C(\sigma_{X})$ such that $\omega(x)\cap Sing(X)=\emptyset$.
Since $\omega(x)$ has no singularities, it follows by Lemma 2.1 that it is hyperbolic of saddle type. In particular, it has local stable and unstable manifolds of size $\varepsilon>0$. So, we obtain that
$W^{u}(z)\pitchfork W^{s}(\gamma_{X})\neq\emptyset$
for some $z\in\omega(x)\subset C(\sigma_{X})$. Thus, since $z$ is chain attainable from $\sigma_{X}$, one has that $\gamma_{X}$ is chain attainable from $\sigma_{X}$.
Case 2: $\omega(x)\cap Sing(X)\neq\emptyset$ for any $x\in C(\sigma_{X})$.
In this case, we denote
The compactness of $C(\sigma_{X})\setminus B(Sing(X|_{C(\sigma_{X})},\delta)$ implies the existence of a positive number $T_{\delta}$ such that if
$z\in C(\sigma_{X})\setminus B(Sing(X|_{C(\sigma_{X})},\delta),$
there is $0<t_{z}\leq T_{\delta}$ such that $X_{t_{z}}(z)\in\mathcal{N}^{in}$ and $X_{(0,t_{z})}(z)\cap\mathcal{N}^{in}=\emptyset$.
By hypothesis on $C(\sigma_{X})$, there is a point $x\in C(\sigma_{X})\setminus W^{s}(Sing(X))$. Moreover, since the positive orbit of $x$ accumulates $Sing(X|_{C(\sigma_{Y})})$, there is a sequence of times $t_{n}\to\infty$ and a singularity $\sigma^{i}\in Sing(X|_{C(X)})$ such that $X_{t_{n}}(x)\to\sigma^{i}$. In particular, the positive orbit of $x$ crosses $\mathcal{N}^{in}$ infinitely many times. Thus, we may assume $x\in\mathcal{N}^{in}_{i}$ for some $i\in\{0,1,\ldots,k\}$. Let us consider a box $B\subset\mathcal{N}^{in}_{i}$ containing $x$ with central volume $0<b<\frac{d^{s}_{Y}}{2}$.
Claim: For any $T>0$, there is $T_{b}>T$, depending on $b$, such that
$X_{[0,T_{b}]}(B)\pitchfork W^{s}(\gamma_{X})\neq\emptyset.$
$K=\inf\left\{\frac{\|X(z)\|}{\|X(w)\|}:z,w\notin B(\sigma_{X},\delta)\right\},$
the sectional-hyperbolicity implies
(5) $)\displaystyle\|\pi_{N_{X_{t}(x)}}(\Phi_{t}(\hat{v}$$)\|$ $=$ $\frac{\|X(x)\|}{\|X(X_{t}(x))\|}\cdot\det(\Phi^{X}_{t}(x)|_{L_{x}})$
$\geq$ $\frac{\|X(x)\|}{\|X(X_{t}(x))\|}\cdot e^{\lambda t}$
$\geq$ $Ke^{\lambda t},$
whenever $t>0$, $X_{[0,t)}(x)\in U$ and $Y_{t}(x)\notin B(\sigma,\delta)$.
Recall by the construction of the cross sections $\mathcal{N}^{in}$ and $\mathcal{N}^{out}$ that if $z\in\mathcal{N}^{in}_{i}$, then $X_{[0,\tau(z))}(z)\in B(\sigma_{X},\delta)\subset U$ and $H_{i}(z)=X_{\tau(z)}(z)\in\mathcal{N}^{out}_{i}$. In this case, define
$\tau_{i}=\inf\{\tau(z):z\in\mathcal{N}^{in}_{i}\}.$
Notice that we can choose $\delta>0$ in a such way that
where $0<\lambda_{u}<\lambda$. Let $\mathcal{N}_{X_{t}(x)}$, $t>\tau(x)$, be a cross-section through $X_{t}(x)$ of radius at least $\eta_{\delta}$. Since $C(\sigma_{Y})$ is not necessarily Lyapunov stable, the set $H^{t}_{x}(B)$ may not be entirely contained in $B(C(\sigma_{X}),2\delta)$ for some $t>\tau_{i}$. Nevertheless, there is a connected component $B_{1}$ of $H_{x}^{t}(B)$ whose central volume is bigger than $b$ and $\overline{B}\subset B(C(\sigma_{X}),2\delta)$. In this case, we set
$r_{1}=\min\{r>0:X_{r}(z)\in\partial B(C(\sigma_{X}),2\delta)\text{ for some }z\in(H^{r}_{x})^{-1}(B_{1})\}.$
Hence, the map
satisfies
$H_{i}(B)\subset Dom(H^{r1}_{x})\textrm{ and }|H^{t}_{x}(B)|_{c}=|H_{x}^{r_{1}}\circ H_{i}(B)|_{c}\geq e^{\lambda^{\prime}(\tau_{i}+r)}b.$
In particular, one has
$|H_{x}^{t}(B)|_{c}\geq e^{\tau_{i}+r_{1}}b>2b.$
By the choice of $\ell$, there is $s\geq 0$ such that $\tau+r_{1}+\cdots+r_{\ell}+s=t^{1}_{x}$. Besides, if $)\hat{w}=\pi_{N_{X_{t_{x}^{1}-s}(x)}}(\Phi_{t^{1}_{x}-s}(\hat{v}$)), we have by (5) and (a) that
$\|\pi_{N_{X_{t_{x}^{1}}(x)}}(\Phi_{s}(\hat{w}))\|\geq\frac{\|X(x)\|}{\|X(X_{t}(x))\|}\cdot e^{\lambda s}\geq e^{\lambda s}>1.$
By repeating the previous process in an inductively way, there is $N\in\mathbb{N}$, and a connected component $C_{N}\subset\mathcal{N}^{in}$ of $H_{x}^{t_{x}^{N}}(B)$ whose central volume is bigger than $Nb>\frac{3}{2}d_{s}^{i}$ for every $i=0,\ldots,k$. Furthermore, we can choose $N$ and $C_{N}$ in a such way that $t_{x}^{N}>T$ and $C_{N}\subset\mathcal{N}_{i}^{in}$. So, we have that $X_{[0,s]}(C_{N})$, $s>0$, is a submanifold of dimension $\dim F$ which intersects $W^{s}(\sigma_{j})$ transversely for some $j\in\{0,1,\ldots,k\}$. Moreover, since the central volume of $C_{N}$ is bigger than $d^{s}_{Y}$, one has by (b) that $W^{s}(\gamma_{j})\pitchfork X_{[0,s_{0}]}(C_{N})\neq\emptyset$ for some $s_{0}>0$. So, the claim follows by taking $T_{b}=t_{x}^{N}+s_{0}$.
To finish the proof, recall that the positive orbit of $x$ crosses $\mathcal{N}^{in}$ infinitely many times and, therefore, there is a sequence of times $t_{n}\to\infty$ and a singularity $\sigma^{i}\in Sing(X|_{C(X)})$ such that $X_{t_{n}}(x)\to x_{0}$ and $x_{0}\in W^{s}_{loc}(\sigma^{i})$. Without loss of generality, we can assume $x_{0},X_{t_{n}}(x)\in\mathcal{N}^{out}_{i}$, for every $n>0$. Fix $\varepsilon$, $T>0$ and fix $N>0$ such that $d(X_{T_{n}}(x),x_{0})\leq\varepsilon$, for every $n\geq N$. Fix $s_{\varepsilon}>T$ satisfying $d(X_{-s_{\varepsilon}}(x_{0}),\sigma^{i})\leq\varepsilon$. Let $x_{n}\in\mathcal{N}^{i}n$ be such that $x_{n}=X_{t_{N}+s}(x)$ and $s>T$. We we consider a box $B$ around $x_{n}$ with diameter $\varepsilon$ and volume $b$, by the claim and the invariance of stable sets, there is $z\in B\cap W^{s}(\gamma_{X})$. Therefore, there exist $s_{z}>T$ and $x_{z}\in\gamma_{X}$ such that $d(X_{s_{z}}(z),x_{z})\leq\varepsilon$. Finally, the set
$\{\sigma^{i},X_{-s_{\varepsilon}}(x_{0}),X_{t_{n}}(x),z,x_{z}\}$
is an $\varepsilon$-$T$-chain from $\sigma^{i}$ to $\gamma_{X}$. Since $\sigma_{X}$ and $\sigma^{i}$ are chain related, the proof is complete. ∎
Finally, we present the next result about the non-trivial dynamical behavior of the chain-recurrence classes $C(\sigma_{Y})$ sufficiently $C^{1}$-close to a chain recurrent class $C(\sigma_{X})$, with $X\in\mathcal{R}$ and such that $C(\sigma_{X})$ is not reduced to either a homoclinic loop or a saddle connection between singularities.
Lemma 3.14.
Let $X\in\mathcal{R}$ and let $C(\sigma_{X})\subset M$ be a non-trivial sectional-hyperbolic chain-recurrent class for $X$ which is not reduced to either a homoclinic loop or a saddle connection between singularities. Then, there is a neighborhood $\mathcal{U}_{2}\subset\mathcal{U}_{1}$ of $X$ such that for every $Y\in\mathcal{U}_{2}$ one has $C(\sigma_{Y})\not\subset W^{s}(Sing(Y))$.
Proof.
@@
By Lemma 3.12, Lemma 3.13 and item (2) of Lemma 3.2 we have that $C(\sigma_{X})$ is a homoclinic class. Take two different periodic orbits $\gamma_{0},\gamma_{1}$ in $C(\sigma_{X})$, and let $d_{\gamma}=d_{H}(\gamma_{0},\gamma_{1})>0$. By continuity of $Y\in\mathcal{U}_{X}\to C(\gamma_{Y})$ there is a neighborhood $\mathcal{U}_{2}\subset\mathcal{U}_{1}$ of $X$ such that
Next, we are ready to prove Theorem 3.1.
Proof of Theorem 3.1.
Let $\mathcal{R}$ be given by Lemma 3.2. We consider the following cases:
Case 1: $\Lambda$ does not contain singularities.
By Lemma 3.4, it is a uniformly hyperbolic homoclinic class. In particular, there exists a neighborhood $\mathcal{U}_{H}$ of $\Lambda$ and a neighborhood $U$ of $\Lambda$ such that the maximal invariant set $H_{Y}$ contained in $U$ for any $Y\in\mathcal{U}_{H}$ is also hyperbolic of saddle type.
Now, by item (3) of Lemma 3.2, there is a neighborhood $\mathcal{U}_{X}\subset\mathcal{U}_{H}$ of $X$ such that the continuation $\Lambda_{Y}$ of $\Lambda$ is contained in $H_{Y}$, so it is also hyperbolic of saddle type. Then, by compactness of $H_{Y}$ and the stable manifold Theorem, there are $\delta,\varepsilon>0$ such that
(7) $W_{\varepsilon}^{s}(z)\pitchfork W_{\varepsilon}^{u}(y)\neq\emptyset,\quad\forall z,y\text{ such that }d(z,y)<\delta.$
Let $x\in\Lambda$. Since $\Lambda$ is a homoclinic class there is a sequence of periodic orbits $\{\gamma_{n}\}_{n\geq 1}\subset\Lambda$ and a sequence of points $p_{n}\in\gamma_{n}$, $n\geq 1$, such that $p_{n}\to x$ when $n\to\infty$. Thus, by $n$ large enough, we can find a neighborhood $\mathcal{U}\subset\mathcal{U}_{X}$ of $X$ such that for any $Y\in\mathcal{U}$, there is a periodic orbit $\gamma_{Y}$ of $Y$ and a point $z\in\Lambda_{Y}$ such that $d(z,y)<\delta$. So, by (7), we have that $z\sim\gamma_{Y}$. This shows that $\gamma_{Y}\subset\Lambda$ by definition of chain class.
Case 2: $\Lambda=C(\sigma_{X})$ for some $\sigma_{X}\in Sing(X)$.
By Lemma 3.12, Lemma 3.13 and item (2) of Lemma 3.2 we have that $C(\sigma_{X})$ is a homoclinic class and there is a neighborhood $\mathcal{U}_{1}$ of $X$ such that the conclusion of Lemma 3.12 holds for every $Y\in\mathcal{U}_{1}$.
On the other hand, by item (3) of Lemma 3.2 and by follow the proof of Lemma 3.13, we deduce that the continuations $Sing(X|_{C(\sigma_{Y})})$ consist entirely of Lorenz-like singularities, for any vector field $Y\in\mathcal{U}_{2}$.
@@
By Lemma 3.14 there is a neighborhood $\mathcal{U}_{2}\subset\mathcal{U}_{1}$ the continuation $C(\sigma_{Y})$, there is a point $x\in C(\sigma_{Y})\setminus W^{s}(Sing(Y))$ for every $Y\in\mathcal{U}_{2}$, so that, by shrinking $\mathcal{U}_{2}$ if it is necessary, we follow Case 1 and Case 2 of the proof of Lemma 3.13 verbatim, replacing $X$ for $Y$, to get a periodic orbit $\gamma_{Y}$ such that $\sigma_{Y}\sim\gamma_{Y}$ and $\gamma_{Y}\sim\sigma_{Y}$. This shows that $\gamma_{Y}\subset C(\sigma_{Y})$ by definition of chain class. ∎
4. Proof of the Main Theorem